# Adversarial Observability and Performance Tradeoffs in Optimal Control

### v1

@@ 4.1 Solution to Problem 1 using an SDP
By the constraints of Problem 1, we have $W=\int_{0}^{\infty}e^{(A+BK)^{\textrm{T}}\tau}C^{\mathrm{T}}Ce^{(A+BK)\tau}\textrm{d}\tau$ and $P=\int_{0}^{\infty}e^{(A+BK)^{\mathrm{T}}\tau}(Q+K^{\mathrm{T}}RK)e^{(A+BK)\tau}\textrm{d}\tau$. Therefore, using the cyclic property of the trace operator, we obtain $\mathrm{tr}(WV)=\mathrm{tr}(CSC^{\mathrm{T}})$ and $\mathrm{tr}(PV)=\mathrm{tr}(S(Q+K^{\mathrm{T}}QK))$, where $S\succ 0$ is the unique solution of the Lyapunov equation
$(A+BK)S+S(A+BK)^{\mathrm{T}}+V=0.$
Using these equivalences, we can reduce the dimension of Problem 1 and conclude it is equivalent to the program
@@
4:     Compute $P,~{}W$ from the equality constraints of Problem 1.
@@ 4.2 Approximate Solution to Problem 2 using Sequential SDP
Toward this end, one challenge entailed in Problem 2 is that its cost function is nonlinear and, most importantly, it decreases rapidly as the matrix $W_{\epsilon}$ approaches singularity. As a result, linearizing this cost can lead to large approximation errors that slow down the convergence of a convex-concave procedure. Nevertheless, by transforming the Lyapunov equation for the observability Gramian into a special form of an algebraic Riccati equation, we can effectively deal with this issue. Specifically, we show Problem 2 can be cast as an equivalent DC program with a linear objective function that involves no inverses of matrix variables.
@@
$-Y_{\epsilon}(A+BK)^{\mathrm{T}}-(A+BK)Y_{\epsilon}-Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}=0,$ (24)
@@
$-Y_{\epsilon}(A+BK+Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I))^{\mathrm{T}}-(A+BK+Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I))Y_{\epsilon}\\ +Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}=0.$
Since $C^{\mathrm{T}}C+\epsilon I\succ 0$, for $Y_{\epsilon}=W_{\epsilon}^{-1}\succ 0$ we have $Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}\succ 0$. By the Lyapunov Theorem, $-(A+BK+Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I))^{\mathrm{T}}$ is strictly stable for $Y_{\epsilon}=W_{\epsilon}^{-1}$, hence $Y_{\epsilon}=W_{\epsilon}^{-1}$ is the stabilizing solution to (24), and hence also the maximal one by Theorem 12.18.4 in [18]. Therefore, for any other solution $\hat{Y}_{\epsilon}\neq W_{\epsilon}^{-1}$ to (24) we have $\hat{Y}_{\epsilon}\prec W_{\epsilon}^{-1}$ and thus $-\mathrm{tr}(\hat{Y}_{\epsilon})>-\mathrm{tr}(W_{\epsilon}^{-1})$. It thus follows that if we substitute the second constraint in Problem 2 with (24) and perform the variable change $Y_{\epsilon}=W_{\epsilon}^{-1}$, the optimal solution will remain the same. Hence, Problem 2 is equivalent to the following optimization problem with linear cost:
@@
$-Y_{\epsilon}(A+BK)^{\mathrm{T}}-(A+BK)Y_{\epsilon}$
$\qquad\qquad\qquad\qquad\qquad\quad-Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}=0,$
@@
$\mathrm{s.t.}\quad$ $\begin{bmatrix}\frac{1}{2}(A{+}BK{+}P)^{\mathrm{T}}(A{+}BK{+}P){+}Q{-}L_{j}(K,P)&K^{\mathrm{T}}\\ \star&-R^{-1}\end{bmatrix}{\preceq}0,$
$\begin{bmatrix}\frac{1}{2}(A{+}BK{+}Y_{\epsilon})(A{+}BK{+}Y_{\epsilon})^{\mathrm{T}}{-}L_{j}^{\prime}(K,Y_{\epsilon})&Y_{\epsilon}\sqrt{C^{\mathrm{T}}C{+}\epsilon I}\\ \star&I\end{bmatrix}{\preceq}0,$
$P\succeq 0,~{}Y_{\epsilon}\succeq 0.$
@@
$\begin{bmatrix}-L_{j}^{\prime}(K,Y_{\epsilon})&Y_{\epsilon}\sqrt{C^{\mathrm{T}}C+\epsilon I}&\frac{1}{\sqrt{2}}(A+BK+Y_{\epsilon})\\ \star&-I&0\\ \star&\star&-I\end{bmatrix}\preceq 0,$
@@
Algorithm 2 requires initialization with sets of matrices $K_{0},P_{0},Y_{\epsilon,0}$ that are feasible for Problem 2. With such an initialization, (26) will also be feasible.
@@
We summarize the convergence properties of Algorithm 2 in the following theorem. This convergence relies fundamentally on the regularization of Problem 2 with the parameter $\epsilon>0$, as shown in the theorem’s proof.
@@
$-Y_{\epsilon}(A+BK)^{\mathrm{T}}-(A+BK)Y_{\epsilon}-Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}\succeq 0.$ (27)
Since the feasible set for $K$ is bounded, $-(A+BK)$ is bounded. Therefore, if $Y_{\epsilon}\succeq 0$ and $\left\lVert Y_{\epsilon}\right\rVert\rightarrow\infty$, then (27) cannot hold because the quadratic term $Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}$ will force the left-hand side of (27) to become negative definite. This is particularly true because $C^{\textrm{T}}C+\epsilon I\succ 0$, following $\epsilon>0$. Hence the feasible set for $Y_{\epsilon}$ is bounded above, hence for any $X\succeq 0$ the set of feasible $Y_{\epsilon}$ satisfying $-\mathrm{tr}(Y_{\epsilon}V^{-1})\leq-\mathrm{tr}(XV^{-1})$ is bounded. Therefore, by [20, 21], Algorithm 2 converges to a stationary point of (23).
@@ 5.1 Intuitive Example
$\dot{x}(t)$ $=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}x(t)+\begin{bmatrix}1\\ 1\end{bmatrix}u(t),$
$y(t)$ $=\begin{bmatrix}1&0\end{bmatrix}x(t).$
@@
Figure 1: Minimum values of Problems 1 and 2 as computed by Algorithm 1 and 2, as well as the corresponding lower bounds calculated in Section 3, for various values of the tradeoff parameter $\lambda$.
@@
Figure 1 shows the tradeoff between performance and adversarial observability. Specifically, it shows the minimum values of Problems 1 and 2 as computed by Algorithm 1 and 2, as well as the corresponding lower bounds we calculated in Section 3, for various values of the tradeoff parameter $\lambda$. We verify that the theoretically derived lower bounds are correct. While they are relatively tight for Problem 1, they become looser for Problem 2 because the inverse of the Gramian tends to be close to singular as $\lambda$ increases, and because the obtained solution is only a local minimum. Finally, we also validate the intuition behind Theorem 3: when minimizing the trace of the observability Gramian, smaller performance tradeoffs yield a larger payoff in adversarial observability. On the other hand, when maximizing the trace of the Gramian’s inverse, the payoff is steeper when the performance tradeoff becomes larger.
@@ 5.2 ADMIRE Aircraft
We perform further simulations on the ADMIRE aircraft [24], which has $n=5$ states and $m=7$ control inputs. We assume that an adversary can measure the first, third, and fifth state, so that
$C=\begin{bmatrix}1&0&0&0&0\\ 0&0&1&0&0\\ 0&0&0&0&1\end{bmatrix},$
while we refer to [24] for the expressions of the matrices $A,B$.
Figure 2: The evolution of the eigenvalues and the trace of the inverse of the observability Gramian of $(A+BK_{j},C)$ during Algorithm 2.
Table 1: Eigenvalues and trace of the observability Gramian under the optimal gain $K^{\star}$ and the gain obtained by Algorithm 1.
$\textrm{tr}(W)$ $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $\lambda_{5}$
Initial $13.8240$ $9.7853$ $3.6252$ $0.3407$ $0.0709$ $0.0018$
Final $11.0108$ $9.3408$ $1.2849$ $0.3313$ $0.0520$ $0.0018$
@@
Figure 2 shows the evolution of the eigenvalues – and the trace of the inverse – of the observability Gramian of $(A+BK_{j},C)$ throughout the execution of Algorithm 2. We observe that the eigenvalue that is affected the most is the minimum one, which is driven very close to zero. This is both an expected and a desirable property. On the one hand, maximizing the trace of the inverse of the observability Gramian is equivalent to making the Gramian singular, which is indeed most easily achieved by making the minimum eigenvalue equal to zero. On the other hand, this outcome means that the resulting closed-loop pair $(A+B\hat{K},C)$ is very close to being unobservable, and hence an adversary who knows $\hat{K}$ and who observes $y(t)$ will not be able to easily reconstruct the full state $x(t)$ of the system. The figure also illustrates a substantial increase in the trace of the observability Gramian, all while maintaining acceptable control performance — as measured by the linear-quadratic cost — with a relatively small compromise, limited to a tradeoff of $\lambda=1$.
@@
Finally, to showcase the adversarial observability and performance tradeoffs provided by Algorithms 1 and 2, we simulate the closed-loop behavior of the system along with a Luenberger observer – used by an adversary to reconstruct the system state – that places observation poles to $-15;-5;-2;-2;-1$. For this state estimation scenario, we assume the sensing components of the adversary are corrupted by deterministic noise composed of $5$ sinusoids, with frequencies between 00 to $\frac{1}{2\pi}$ and magnitudes equal to $0.01$. Figure 3 shows the incurred closed-loop performance cost along with the estimation errors of the adversary under a) the nominal optimal controller; b) the controller obtained by Algorithm 1; and c) the controller obtained by Algorithm 2. We observe that the controller for Algorithm 1 magnifies adversarial estimation errors across all states with almost negligible loss of performance, while Algorithm 1 increases the estimation error of the fourth state to completely unreliable levels with a little more performance tradeoff. This pattern also aligns with the interpretation of the metrics (4): on the one hand, the trace of the observability Gramian captures estimation capabilities across all directions in the state space; on the other hand, the trace of the observability Gramian inverse is skewed toward the least observable state which, in this example, corresponds to the fourth state of the aircraft.
@@ 6 Conclusion
We study the problem of feedback control that minimizes the observability of a set of sensors deployed by an adversary, while conforming to strict closed-loop performance constraints. We quantify observability using metrics related to the trace of the observability Gramian and its inverse. For both of these, we establish theoretical lower bounds on their optimal values as a function of the performance tradeoff parameter. Finally, we optimize the trace of the observability Gramian through an SDP, while, for the trace of the Gramian inverse, we perform this optimization approximately using sequential SDP.

### v2

@@ 4.1 Solution to Problem 1 using an SDP
By the constraints of Problem 1, we have $W=\int_{0}^{\infty}e^{(A+BK)^{\textrm{T}}\tau}C^{\mathrm{T}}Ce^{(A+BK)\tau}\textrm{d}\tau$ and $P=\int_{0}^{\infty}e^{(A+BK)^{\mathrm{T}}\tau}(Q+K^{\mathrm{T}}RK)e^{(A+BK)\tau}\textrm{d}\tau$. Therefore, using the cyclic property of the trace operator, we obtain $\mathrm{tr}(WV)=\mathrm{tr}(CSC^{\mathrm{T}})$ and $\mathrm{tr}(PV)=\mathrm{tr}(S(Q+K^{\mathrm{T}}QK))$, where $S\succ 0$ is the unique solution of the Lyapunov equation $(A+BK)S+S(A+BK)^{\mathrm{T}}+V=0.$ Using these equivalences, we can reduce the dimension of Problem 1 and conclude it is equivalent to the program
@@ 4.2 Approximate Solution to Problem 2 using Sequential SDP
4:  Compute P,WP,~W from the equality constraints of Problem 1.
@@
Toward this end, one challenge entailed in Problem 2 is that its cost function is nonlinear and, most importantly, it decreases rapidly as the matrix $W_{\epsilon}$ approaches singularity. As a result, linearizing this cost can lead to large approximation errors that slow down the convergence of a convex-concave procedure. Nevertheless, by transforming the Lyapunov equation for the observability Gramian into a special form of an algebraic Riccati equation, we can effectively deal with this issue. Specifically, we show Problem 2 can be cast as an equivalent difference of convex functions (DC) program with a linear objective function that involves no inverses of matrix variables.
@@
$Y_{\epsilon}(A+BK)^{\mathrm{T}}-(A+BK)Y_{\epsilon}-Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}=0,$ (25)
@@
$\hskip-8.53581pt-Y_{\epsilon}(A+BK+Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I))^{\mathrm{T}}-(A+BK+Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I))Y_{\epsilon}\\ +Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}=0.$
Since $C^{\mathrm{T}}C+\epsilon I\succ 0$, for $Y_{\epsilon}=W_{\epsilon}^{-1}\succ 0$ we have $Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}\succ 0$. By the Lyapunov Theorem, $(A+BK+Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I))^{\mathrm{T}}$ is strictly stable for $Y_{\epsilon}=W_{\epsilon}^{-1}$, hence $Y_{\epsilon}=W_{\epsilon}^{-1}$ is the stabilizing solution to (25), and hence also the maximal one by Theorem 12.18.4 in [3]. Therefore, for any other solution $\hat{Y}_{\epsilon}\neq W_{\epsilon}^{-1}$ to (25) we have $\hat{Y}_{\epsilon}\prec W_{\epsilon}^{-1}$ and thus $\mathrm{tr}(\hat{Y}_{\epsilon}V^{-1})>-\mathrm{tr}(W_{\epsilon}^{-1}V^{-1})$. It thus follows that if we substitute the second constraint in Problem 2 with (25) and perform the variable change $Y_{\epsilon}=W_{\epsilon}^{-1}$, the optimal solution will remain the same. Hence, Problem 2 is equivalent to the following optimization problem with linear cost:
@@
$-Y_{\epsilon}(A+BK)^{\mathrm{T}}-(A+BK)Y_{\epsilon}-Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}=0,$
@@
$\mathrm{s.t.}\quad$ $\hskip-5.69054pt\begin{bmatrix}\frac{1}{2}(A{+}BK{+}P)^{\mathrm{T}}(A{+}BK{+}P){+}Q{-}L_{j}(K,P)&K^{\mathrm{T}}\\ \star&-R^{-1}\end{bmatrix}{\preceq}0,$
$\hskip-25.60747pt\begin{bmatrix}\frac{1}{2}(A{+}BK{+}Y_{\epsilon})(A{+}BK{+}Y_{\epsilon})^{\mathrm{T}}{-}L_{j}^{\prime}(K,Y_{\epsilon})&Y_{\epsilon}\sqrt{C^{\mathrm{T}}C{+}\epsilon I}\\ \star&I\end{bmatrix}{\preceq}0,$
@@
$\hskip-17.07164pt\begin{bmatrix}-L_{j}^{\prime}(K,Y_{\epsilon})&Y_{\epsilon}\sqrt{C^{\mathrm{T}}C+\epsilon I}&\frac{1}{\sqrt{2}}(A+BK+Y_{\epsilon})\\ \star&-I&0\\ \star&\star&-I\end{bmatrix}\preceq 0,$
@@
Algorithm 2 requires initialization with sets of matrices $K_{0},P_{0},Y_{\epsilon,0}$ that are feasible for Problem 2. One such initialization is obtained by selecting $K_{0}$ as the linear-quadratic gain corresponding to $\lambda=0$. The matrices $P_{0}$ and $Y_{\epsilon,0}$ are then obtained by solving the first and second Lyapunov equations of Problem 2 under $K=K_{0}$, respectively, with $Y_{\epsilon,0}$ computed as the inverse of the resulting $W_{\epsilon}$. With such an initialization, (27) will also be feasible.
@@
We summarize the convergence properties of Algorithm 2 in the following theorem. This convergence relies fundamentally on the regularization of Problem 2 with the parameter $\epsilon>0$.
@@
$Y_{\epsilon}(A+BK)^{\mathrm{T}}-(A+BK)Y_{\epsilon}-Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}\succeq 0.$ (28)
Since the feasible set for $K$ is bounded, (A+BK)-(A+BK) is bounded. Therefore, if $Y_{\epsilon}\succeq 0$ and $\left\lVert Y_{\epsilon}\right\rVert\rightarrow\infty$, then (28) cannot hold because the quadratic term $Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}$ will force the left-hand side of (28) to become negative definite. This is particularly true because $C^{\textrm{T}}C+\epsilon I\succ 0$, following $\epsilon>0$. Hence the feasible set for $Y_{\epsilon}$ is bounded above, hence for any $X\succeq 0$ the set of feasible $Y_{\epsilon}$ satisfying $\mathrm{tr}(Y_{\epsilon}V^{-1})\leq-\mathrm{tr}(XV^{-1})$ is bounded. Therefore, by [18, 17], Algorithm 2 converges to a stationary point of (24).
@@ 5.1 Intuitive Example
$\dot{x}(t)$ $=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}x(t)+\begin{bmatrix}1\\ 1\end{bmatrix}u(t),\quad y(t)=\begin{bmatrix}1&0\end{bmatrix}x(t).$
@@
Figure 2: Minimum values of Problems 1 and 2 as computed by Algorithm 1 and 2, as well as the corresponding lower bounds calculated in Section 3, for various values of the trade-off parameter $\lambda$.
@@
Figure 2 shows the trade-off between performance and adversarial observability. Specifically, it shows the minimum values of Problems 1 and 2 as computed by Algorithm 1 and 2, as well as the corresponding lower bounds we calculated in Section 3, for various values of the trade-off parameter $\lambda$. We verify that the theoretically derived lower bounds are correct. While they are relatively tight for Problem 1, they become looser for Problem 2 because the inverse of the Gramian tends to be close to singular as $\lambda$ increases, and because the obtained solution is only a local minimum. Finally, we also validate the intuition behind Theorem 3: when minimizing the trace of the observability Gramian, smaller performance trade-offs yield a larger payoff in adversarial observability. On the other hand, when maximizing the trace of the Gramian’s inverse, the payoff is steeper when the performance trade-off becomes larger.
@@ 5.2 ADMIRE Aircraft
We perform further simulations on the ADMIRE aircraft [15], which has $n=5$ states and $m=7$ control inputs. We assume that an adversary can measure the first, third, and fifth state, while we refer to [15] for the expressions of the matrices $A,B$.
Figure 3: Evolution of the eigenvalues $\lambda_{i}(W)$, $i=1,\ldots,5$, of the observability Gramian $W$ and of the trace of its inverse, $\mathrm{tr}(W^{-1})$, for $(A+BK_{j},C)$ during Algorithm 2.
Table 1: Eigenvalues under the nominal optimal gain $K^{\star}$ and the gain obtained by Algorithms 1-2, and the corresponding costs.
$\textrm{tr}(W)$ $\textrm{tr}(W^{-1})$ $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $\lambda_{5}$
Nom. $13.82$ $558$ $9.79$ $3.63$ $0.341$ $0.0709$ $0.00180$
Alg. 1 $11.01$ $598$ $9.34$ $1.28$ $0.331$ $0.0520$ $0.00180$
Alg. 2 $13.80$ $31504$ $9.78$ $3.60$ $0.351$ $0.0707$ $0.00003$
@@
Figure 3 shows the evolution of the eigenvalues – and the trace of the inverse – of the observability Gramian of $(A+BK_{j},C)$ throughout the execution of Algorithm 2. We observe that the eigenvalue that is affected the most is the minimum one, which is driven very close to zero (see also Table 1). This is both an expected and a desirable property. On the one hand, maximizing the trace of the inverse of the observability Gramian is equivalent to making the Gramian singular, which is indeed most easily achieved by making the minimum eigenvalue equal to zero. On the other hand, this outcome means that the resulting closed-loop pair $(A+B\hat{K},C)$ is very close to being unobservable, and hence an adversary who knows $\hat{K}$ and who observes $y(t)$ will not be able to easily reconstruct the full state $x(t)$ of the system. The figure also illustrates a substantial increase in the trace of the observability Gramian, all while maintaining acceptable control performance — as measured by the linear-quadratic cost — with a relatively small compromise, limited to a trade-off of $\lambda=1$.
@@
Finally, to showcase the adversarial observability and performance trade-offs provided by Algorithms 1 and 2, we simulate the closed-loop behavior of the system along with a Luenberger observer – used by an adversary to reconstruct the system state – that places observation poles to 15;5;2;2;1-15;-5;-2;-2;-1. For this state estimation scenario, we assume the sensing components of the adversary are corrupted by deterministic noise composed of $5$ sinusoids, with frequencies between 0 to $\frac{1}{2\pi}$ and magnitudes equal to $0.01$. Figure 4 shows the incurred closed-loop performance cost along with the estimation errors of the adversary under a) the nominal optimal controller; b) the controller obtained by Algorithm 1; and c) the controller obtained by Algorithm 2. We observe that the controller for Algorithm 1 magnifies adversarial estimation errors across all states with almost negligible loss of performance, while Algorithm 1 increases the estimation error of the fourth state to completely unreliable levels with a little more performance trade-off. This pattern also aligns with the interpretation of the metrics (4): on the one hand, the trace of the Gramian captures estimation capabilities across all directions in the state space; on the other hand, the trace of the Gramian inverse is skewed toward the least observable state which, in this example, corresponds to the fourth state of the aircraft.
@@ 6 Conclusion
We study the problem of feedback control that minimizes the observability of a set of sensors deployed by an adversary, while conforming to strict closed-loop performance constraints. We quantify observability using metrics related to the trace of the observability Gramian and its inverse. For both of these, we establish theoretical lower bounds on their optimal values as a function of the performance trade-off parameter. Finally, we optimize the trace of the observability Gramian through an SDP, while, for the trace of the Gramian inverse, we perform this optimization approximately using sequential SDP.